# Assist Research

### v1

@@ 5.2 ExperimentDesign
Table 2: Various LLMs’ performances on the 100 instances of ExpDesign. The explanation generation is based on the oracle experiments to prevent error propagation. “Copy Input” is a random baseline: for experiment design, randomly select 5 sentences from the input paper; for experiment explanation, directly copy each experiment idea.
@@
Methods S-F1 S-Precision S-Recall S-Match ROUGE-L ROUGE-1
Copy Input 21.13 17.94 26.76 40.32 22.06 25.28
@@
OLMo-7B (Groeneveld et al., 2024) 33.94 37.25 31.79 45.78 26.30 30.38
Falcon-40B (Almazrouei et al., 2023) 17.87 21.78 15.35 17.03 12.10 12.72
Gemma 2-27B (Gemma Team, 2024) 34.33 39.71 30.51 42.77 26.20 29.63
Mistral-7B (Jiang et al., 2023) 37.62 43.09 34.19 50.18 30.20 34.69
Mixtral-8x22B-MoE (Jiang et al., 2024) 42.21 50.13 36.82 49.07 29.96 34.53
Llama 3.1-70B (MetaAI, 2024) 40.57 48.43 35.43 50.05 29.33 34.11
Qwen 2.5-72B (Qwen Team, 2024) 43.24 51.73 37.55 51.12 29.46 34.68
@@
Gemini 1.5 Pro (Anil et al., 2023) 51.87 50.77 53.37 52.87 28.52 33.80
Claude 3.5 sonnet (Anthropic, 2024a) 48.74 46.49 51.53 53.03 18.75 26.15
GPT-4 (OpenAI et al., 2023) 43.89 42.34 45.82 55.03 22.82 30.01
GPT-4o (OpenAI, 2024a) 53.00 51.24 55.12 54.79 27.54 34.31
o1-preview (OpenAI, 2024b) 46.67 45.04 48.70 58.55 29.11 36.70
@@
Table 2 shows the main results. For the experiment design, the closed-source LLMs generally outperform open-source LLMs, and both closed-/open-source LLMs are superior to the “Copy Input” baseline (except the Falcon). Despite the higher S-Precision, the open-source LLMs are seriously deficient in S-Recall compared with closed-source LLMs ($\sim$10$\downarrow$). We find that closed-source LLMs are more creative in experiment design and tend to generate more experiment ideas than open-source LLMs (though most of the experiment ideas are trivial), leading to excellent S-Recall. As for the experiment explanation, the S-Match scores of closed-source LLMs still surpass the open-source LLMs, while the score difference is not significant. Furthermore, we find the negative correlation between S-Match and the ROUGE, where the ROUGE scores of closed-source LLMs are broadly inferior. We find that the open-source LLMs often try to copy the terms or phrases from the given experiment, or even simply paraphrase the experiment instead of explaining, which results in a high superficial overlap with the ground-truth explanation. This observation highlights the importance of adopting the proposed S-Match to avoid evaluation bias of traditional generation metrics.
$\mathcal{Q}_{1}$: Can self-contained experiment design enhance the experiment explanation?
@@
$\mathcal{Q}_{2}$: Do human evaluation results align with automatic metrics for explanation?
@@
$\mathcal{Q}_{3}$: Do more contexts boost performance?
We also investigate the impact of input context length for ExpDesign. As shown in Figure 4, we scale up the input pre-experiment context length from 0.1k to 10k tokens (10k tokens is the maximum paper context length in the dataset). For the experiment planning, more input context does improve the performance of different LLMs, while this benefit stops after exceeding 5k tokens, which is similar to EqInfer’s scaling results — after the necessary information has been covered, scaling more up doesn’t boost the performance. Meanwhile, the results of the experiment explanation demonstrate that explaining motivations almost doesn’t require any paper context, i.e., the LLMs solely rely on the given experiments. However, we do not expect this because we hope LLMs can explain the motivation based on a thorough understanding of the paper, just like how human experts do. Hence, there is still a considerable gap between the LLMs and humans in terms of grasping research motivations.
$\mathcal{Q}_{4}$: Does multi-modal input boost performance?
Experiment Design Experiment Explanation
Models S-F1 S-Precision S-Recall S-Match ROUGE-L ROUGE-1
GPT-4o 53.00 51.24 55.12 58.54 29.25 35.50
w/ figures 50.11 48.94 51.59 58.53 27.87 34.30
GPT-4 43.89 42.34 45.82 56.95 25.98 33.37
w/ figures 43.54 42.56 44.85 55.03 22.82 30.01
InternVL2-26B 40.52 48.95 35.20 50.03 29.13 34.26
w/ figures 38.83 46.91 33.70 50.29 29.29 34.06
Intuitively, besides the text, when designing experiments for a given research topic, the figures can provide rich supplementary information, such as an algorithm illustration that can help better understand this research topic and underlying motivations. Hence, we test the performance of different LMMs (Large Multimodal Models), including GPT4-o, GPT-4, and InternVL2 (Chen et al., 2024a). Table 5 shows the ablation results on the figure data. To our surprise, the figure data doesn’t improve the LMMs’ results in this task, even harming the performances. This might be due to the low informativeness of the figures, as figures usually consume more input tokens but act only as supplementary information to the text, indicating future work on developing LMMs that can effectively leverage the scientific figures.
5.3 PaperWeakness
Settings.
Intuitively, the full paper content is necessary for providing feedback to that paper. Therefore, instead of setting a maximum input length, in Weakness, we try to feed all the paper context into the LLMs. As the input length of Weakness is extremely long (see Table 13), we adopt a “split-combine” method — we first split the whole paper into several smaller pieces and let LLMs predict the weaknesses of each piece separately; after that, we combine all pieces’ weaknesses as a final complete prediction. In practice, for the length of each small piece, we set 2,000 and 3,000 words for open- and closed-source LLMs, respectively. Additionally, in this task, we also examine the performance of a recent agent framework, namely AI-SCI (Lu et al., 2024), which enhances GPT-4o’s paper review ability by leveraging advanced prompting techniques, e.g., self-reflection (Shinn et al., 2024) and response ensembling (Wang et al., 2023). [^4] We don’t run AI-SCI on ExpDesign, because AI-SCI takes model-generated ideas as the inputs, which are incompatible with our task setting.
Main results.
@@ 5.3 PaperWeakness
Methods SN-F1 (%) SN-Precision (%) SN-Recall (%) ITF-IDF ($\uparrow$)
@@
Table 6 shows the main results, where the closed-source LLMs’ overall performances are generally superior to the results of open-source LLMs. Similarly, closed-source LLMs are particularly excellent in SN-Recall because of more generated weaknesses. However, there is still a considerable gap in the weakness diversity between the LLMs and human experts. [^5] Note that the human’s ITF-IDF score in Table 6 can be slightly underestimated. This is because we keep the repeated weaknesses in the human review, which affects the human review’s informativeness (lower ITF) but is useful when calculating the SN-Recall for LLMs. Compared with human review, most LLM-generated weaknesses are vague and lack the necessary knowledge about some frontier research works. Surprisingly, AI-SCI performs worse than backbone GPT-4o, especially on ITF-IDF, which suggests the challenge of Weakness, i.e., simply adopting popular prompting techniques cannot well address this task.
Models Input Context Processing Window Size (in words) SN-F1 SN-Precision SN-Recall ITF-IDF
GPT-4-Turbo split-combine 3,000 47.66 42.15 55.19 5.31
no-split 3,000 45.80 43.66 48.39 5.58
no-split 20,000 44.99 42.64 47.82 5.58
GPT-4o split-combine 3,000 47.73 42.09 55.48 5.95
no-split 3,000 45.74 43.45 48.54 5.92
no-split 20,000 45.47 42.97 48.51 6.02
AI-SCI split-combine 3,000 45.05 40.02 51.91 2.23
no-split 3,000 42.56 40.90 44.65 2.53
no-split 20,000 42.53 40.75 44.78 2.58
@@
Ideally, if the LLM has a sufficient context window size, it is unnecessary to split the input papers for separate processing. Consequently, in this paragraph, we utilize the LLMs accepting long context input to compare “split-combine” with “no-split”, i.e., letting LLMs write weaknesses by giving the full paper. In practice, we set the maximum number of input words to 20k, which ensures $\geq$95papers in the Weakness can be fully processed. As shown in Table 7, compared with giving the full paper contexts, split-combine generally brings about superior performances. During manual checking, we find that, when full paper is available, LLMs frequently neglect some important sections and omit weaknesses accordingly, while split-combine ensures that the LLMs can carefully brainstorm weaknesses within each smaller piece. Surprisingly, the LLMs’ performances with full paper context can be even worse than just remaining the first 3,000 words. This implies that even the current powerful long-context LLMs still fall short when processing long scientific documents.
@@
Models SN-F1 SN-Precision SN-Recall ITF-IDF
GPT-4o 47.73 42.09 55.48 5.95
w/ tables 46.76 41.32 54.17 5.53
w/ figures 46.62 41.20 54.04 5.48
w/ tables & figures 46.58 41.17 53.98 5.36
InternVL2-26B 41.91 41.02 43.28 1.48
w/ tables 40.55 40.37 42.91 1.46
w/ figures 42.88 42.10 43.76 1.46
w/ tables & figures 42.44 42.00 43.31 1.44
@@ 6 Conclusion
In this work, we propose AAAR-1.0, a novel benchmark targeting a comprehensive evaluation of the current LLMs’ AI research capacity. AAAR-1.0 consists of distinct expertise-intensive tasks along with the curated evaluation metrics, and collect high-quality data by employing senior AI researchers. Extensive experiments highlight the challenges and values of AAAR-1.0.
Limitations
We shed light on two limitations of this work: i) As we gather data from open-source platforms such as arXiv and OpenReview, there is a possibility that current or future LLMs may be trained on the same source data utilized in our benchmark. This situation could influence the fairness of LLM comparisons and the conclusions drawn from this paper. While we acknowledge this potential data leakage, we maintain that our work can provide valuable insights and serve as an upper bound for some LLMs, particularly if they are indeed pretrained on those scientific papers. At present, this research can inform the design of current LLMs and agents, enhancing the community’s understanding of the strengths and limitations of using LLMs in scientific research. ii) Meanwhile, due to the high consumption of employing senior researchers in conducting data annotation, the data size for some tasks, such as ExpDesign, is relatively small. This might introduce variability in LLM performance, even with repeated runs. As this work marks the beginning of a series, we plan to release larger-scale datasets that will cover more diverse research tasks in AAAR-2.0.
Ethics Statement
Our study highlights that LLMs can be used to assist humans in AI research. This doesn’t mean we encourage the use of AI to replace human researchers. In contrast, we emphasize that the main responsibility for conducting scientific research should always lie with humans to prevent any societal risks, while the LLMs are only tools for making human research more efficient. To this end, we commit to careful distribution of the data collected in our research, ensuring it serves strictly for research purposes. Our goal is to mitigate the risks while maximizing the benefits offered by LLMs.
Acknowledgments
We would like to thank arXiv and OpenReview for releasing the paper source packages and the review comments. The authors would also like to thank Ibraheem Moosa and Sarkar Snigdha Sarathi Das for assisting in the data collection.
@@ Appendix A Data Statistics and Diversity
We provide the detailed data statistics of three datasets in our benchmark, as shown in Table 11, 12, and 13. We use the NLTK package [^7] https://www.nltk.org/ to tokenize words and count the length. When calculating the length of equations, we use the pylatexenc tool [^8] https://github.com/phfaist/pylatexenc to simplify the equations first.
@@
# of classification instances 1,049
@@ Appendix C More Experiment Results and Details
C.1 Human Evaluation on LLM-Generated Explanation
@@
We ask 5 annotators to evaluate the LLM-generated explanations. Specifically, each of them is assigned 4 or 5 papers, along with the corresponding experiment lists. For each paper, the annotator is given 5 different models’ outputs (model names are anonymized), and the annotator has to decide if each LLM-generated explanation is acceptable according to the experiment. We show the human evaluation results in Table 4,

### v2

@@ 5.2 ExperimentDesign
Table 2: Various LLMs’ performances on the 100 instances of ExpDesign. The explanation generation is based on the oracle experiments to prevent error propagation. “Copy Input” directly copies each experiment idea as the explanation.
@@
Methods En-F1 En-Precision En-Recall S-Match ROUGE-L ROUGE-1
Copy Input 40.32 22.06 25.28
@@
OLMo-7B (Groeneveld et al., 2024) 14.80 17.50 19.80 45.78 26.30 30.38
Mistral-7B (Jiang et al., 2023) 18.96 24.83 21.38 50.18 30.20 34.69
Mixtral-8x22B-MoE (Jiang et al., 2024) 23.16 24.45 30.57 49.07 29.96 34.53
Llama 3.1-70B (MetaAI, 2024) 22.92 23.10 29.76 50.05 29.33 34.11
Qwen 2.5-72B (Qwen Team, 2024) 24.28 22.48 34.44 51.12 29.46 34.68
@@
Gemini 1.5 Pro (Anil et al., 2023) 27.25 28.66 34.92 52.87 28.52 33.80
Claude 3.5 sonnet (Anthropic, 2024a) 27.99 24.48 42.09 53.03 18.75 26.15
GPT-4o (OpenAI, 2024a) 25.03 22.25 36.59 54.79 27.54 34.31
o1-preview (OpenAI, 2024b) 30.13 28.13 38.59 58.55 29.11 36.70
@@
Table 2 shows the main results. For the experiment design, the closed-source LLMs generally outperform open-source LLMs. However, the score values of all LLMs are relatively low (20$\sim$30%), implying the LLMs consistently miss ground-truth experiments from the origin paper (low recall), and they tend to generate more novel experiments that didn’t show in the origin paper (low precision). As for the experiment explanation, the S-Match scores of closed-source LLMs still surpass the open-source LLMs. Furthermore, there is a negative correlation between S-Match and ROUGE score, where the ROUGE scores of closed-source LLMs are broadly inferior. We find that the open-source LLMs often try to copy the terms or phrases from the given experiment, or even simply paraphrase the experiment instead of explaining, which results in a high superficial overlap with the ground-truth explanation. This observation highlights the importance of adopting the proposed S-Match to avoid evaluation bias of traditional generation metrics.
$\mathcal{Q}_{1}$: What is the quality of the model-generated novel experiments?
The low En-Precision of LLMs in Table 2 indicates the creativity of LLMs in generating novel experiments. We then randomly sample 15 papers from the ExpDesign and ask 3 experts to manually review the model-generated novel experiments. Specifically, we ask the experts to judge the necessity of the novel experiments, where we set three necessity levels: “A” indicates the experiment is necessary/mandatory to support the main claim, “B” represents optional/supplementary experiments, and “C” for those unrelated experiments (see Appendix C.2 for evaluation details). Table 3 shows the necessity scores of the three strongest LLMs. We find that LLMs consistently generate a lot of novel experiments, especially the Claude; though most of them are optional, even fancy/unrelated experiments, there are still a considerable amount of necessary experiments generated, e.g., the results of o1. We further find that some novel experiments can be regarded as useful supplementary analyses w.r.t. the human experiments. Table 13 shows examples of model-suggested experiments.
Table 3: The human evaluation results on the novel experiments suggested by LLMs. “A”, “B”, and “C” represent the different quality level (i.e., necessity); “A” is the best level.
Models # of novel EXP Necessity (%)
A B
Gemini 1.5 Pro 59 30.59 45.76
Claude 3.5 sonnet 112 21.78 50.00
o1-preview 71 35.84 36.61
$\mathcal{Q}_{2}$: Can self-contained experiment design enhance the experiment explanation?
@@
$\mathcal{Q}_{3}$: Do human evaluation results align with automatic metrics for explanation?
@@
Methods S-F1 (%) S-Precision (%) S-Recall (%) ITF-IDF ($\uparrow$)
@@
$\mathcal{Q}_{4}$: Do more contexts boost performance?
We also investigate the impact of input context length for ExpDesign. As shown in Figure 5, we scale up the input pre-experiment context length from 0.1k to 10k tokens (10k is the length of the longest paper). For the experiment design, more input context does improve the performance of different LLMs, while this benefit stops after exceeding 8k tokens, which means that after the necessary information has been covered, scaling context becomes inefficient. Meanwhile, the explanation generation results reveal that LLMs primarily depend on given experiments rather than paper context to explain motivations. However, we do not expect this as we hope LLMs can explain the motivation based on a thorough understanding of the paper, just like how human experts do. Hence, there is still a considerable gap between the LLMs and humans in terms of grasping research motivations.
$\mathcal{Q}_{5}$: Does multi-modal input boost performance?
Intuitively, besides the text, when designing experiments for a given research topic, the figures can provide rich supplementary information, such as an algorithm illustration that can help better understand this research topic and underlying motivations. Hence, we test the performance of different LMMs (Large Multimodal Models), including GPT4-o and InternVL2 (Chen et al., 2024b). Table 14 shows the ablation results on the figure data. To our surprise, the figure data doesn’t improve the LMMs’ results in this task, even harming the performances. This might be due to the low informativeness of the figures, as figures usually consume more input tokens but act only as supplementary information to the text, indicating future work on developing LMMs that can effectively leverage the scientific figures.
5.3 PaperWeakness
Settings.
Intuitively, full paper content is necessary for paper reviewing. Therefore, instead of setting a maximum input length, in Weakness, we try to utilize the whole paper. As the input length of Weakness is extremely long (see Table 11), we adopt a “split-combine” method — we first split the whole paper into smaller pieces and let LLMs predict the weaknesses of each piece separately; after that, we merge all pieces’ weaknesses as a final prediction. For the length of each small piece, we set 2,000 and 3,000 words for open- and closed-source LLMs, respectively. Additionally, in this task, we also examine the performance of AI-SCI (Lu et al., 2024), which enhances LLMs’ paper review ability by leveraging advanced prompting techniques, e.g., self-reflection (Shinn et al., 2024) and response ensembling (Wang et al., 2023). [^6] We don’t run AI-SCI on ExpDesign, because AI-SCI takes model-generated ideas as the inputs, which are incompatible with our task setting.
Main results.
Table 6 shows the main results, where the closed-source LLMs’ overall performances are generally superior to the results of open-source LLMs. Similarly, closed-source LLMs are particularly excellent in S-Recall because of more generated weaknesses. However, there is still a considerable gap in the weakness diversity between the LLMs and human experts. [^7] The human’s ITF-IDF score in Table 6 can be slightly underestimated. This is because we keep the repeated weaknesses in the human review, which affects the human review’s informativeness (lower ITF) but is useful when calculating the S-Recall for LLMs. Compared with human review, most LLM-generated weaknesses are vague and lack the necessary knowledge about some frontier research works. Surprisingly, AI-SCI performs worse than backbone GPT-4o, especially on ITF-IDF, which suggests the challenge of Weakness, i.e., simply adopting popular prompting techniques cannot well address this task.
@@ 5.3 PaperWeakness
Ideally, if the LLM has a sufficient context window size, splitting the input papers for separate processing is unnecessary. Consequently, in this paragraph, we utilize the LLMs accepting long context input to compare “split-combine” with “no-split”, i.e., letting LLMs write weaknesses by giving the full paper. In practice, we set the maximum number of input words to 20k, which ensures $\geq$95papers in the Weakness can be fully processed. As shown in Table 12, compared with giving the full paper contexts, split-combine generally brings about superior performances. During manual checking, we find that, when full paper is available, LLMs frequently neglect some important sections and omit weaknesses accordingly, while split-combine ensures that the LLMs can carefully brainstorm weaknesses within each smaller piece. Surprisingly, the LLMs’ performances with full paper context can be even worse than just remaining the first 3,000 words. This implies that even the current powerful long-context LLMs still fall short when processing long scientific documents.
@@ 6 Conclusion
In this work, we propose AAAR-1.0, a novel benchmark targeting a comprehensive evaluation of the current LLMs’ capacity in assisting AI research. AAAR-1.0 consists of distinct expertise-intensive tasks along with the curated evaluation metrics. We collect high-quality data by employing senior AI researchers and conducting strict data examinations. Extensive experiments highlight the challenges and values of AAAR-1.0.
Impact Statement
Our study explores whether LLMs can assist human researchers in AI research. We do not advocate for AI replacing human researchers. Instead, we stress that the primary responsibility for scientific research should remain with humans to prevent societal risks, with LLMs serving as tools to enhance research efficiency. Specifically, our work analyzes the strengths and weaknesses of LLMs to ensure researchers remain judicious in their use of these tools. Our goal is to mitigate risks while maximizing the benefits offered by LLMs. We are committed to the careful distribution of data collected in our research, ensuring it is used solely for research purposes.
@@ Appendix A Data Statistics and Diversity
We provide the detailed data statistics of three datasets in our benchmark, as shown in Table 9, 10, and 11. We use the NLTK package [^9] https://www.nltk.org/ to tokenize words and count the length. When calculating the length of equations, we use the pylatexenc tool [^10] https://github.com/phfaist/pylatexenc to simplify the equations first.
@@
# of positive equations 1,049
# of negative equations 3,147
@@ Appendix C More Experiment Results and Details
C.1 Input Context Scaling Investigation
Figure 4, Figure 5, and Table 12 show the context scaling results of EqInfer, ExpDesign, and Weakness.
Models Input Context Processing Window Size (in words) S-F1 S-Precision S-Recall ITF-IDF
GPT-4o split-combine 3,000 47.73 42.09 55.48 5.95
no-split 3,000 45.74 43.45 48.54 5.92
no-split 20,000 45.47 42.97 48.51 6.02
AI-SCI split-combine 3,000 45.05 40.02 51.91 2.23
no-split 3,000 42.56 40.90 44.65 2.53
no-split 20,000 42.53 40.75 44.78 2.58
Figure 4: The input context length scaling trend on the EqInfer task.
C.2 Human Evaluation on LLM-Generated Novel Experiments
Figure 6: The human guideline for evaluating the LLM-generated novel experiments.
Figure 6 illustrates the evaluation guideline for novel experiments generated by LLMs. We ask 3 senior PhD students to evaluate each paper; that is, if the first two annotators disagree with each other, a third annotator will make a final decision. Table 13 presents several human evaluation cases.
Table 13: Examples of human evaluation on the model-generated novel experiments.
Paper Title Original Experiments (by human) Novel Experiment (by LLMs) Rating
WiCE: Real-World Entailment for Claims in Wikipedia 1. Analysis in Verification Problem Distribution: This paper should provide detailed analysis and statistics about the verification problems in the proposed dataset. 2. Off-the-shelf entailment classification performance: The authors should provide entailment classification performance of existing models on the proposed dataset without fine-tuning. 3. Human Performance: The authors should show human performance on the proposed dataset. 4. Performance of fine-tuned models: The authors should provide the performance of models fine-tuned on the proposed dataset. 5. Performance on the evidence retrieval task: The authors should show the performance on the evidence retrieval task, which is a sub-task of the proposed dataset. 6. Performance of LLMs: The authors should provide the performance of LLMs on the proposed dataset. 7. Retrieval+Entailment: Authors should provide experiments on a framework of retrieving evidence sentences and evaluate entailment by using the retrieved sentences. 8. Analysis of Claim-Split on Downstream Tasks: The authors should analyze how claim-split, the proposed method, is effective on tasks other than the proposed dataset. Assess model performance on WiCE without fine-tuning to test domain generalization from traditional NLI datasets. A
MetaMath: Bootstrap Your Own Mathematical Questions for Large Language Models 1. Results of multiple LLMs on popular math datasets: The authors should show the performance of multiple LLMs fine-tuned on their dataset on popular math datasets. 2. Performance on open-source models with different sizes: The authors should show the performance of models with different sizes trained on the proposed dataset. 3. Comparison to SOTA closed-source models: The authors show compare the performance of open-source models trained on the proposed dataset and strong close-source models. 4. Evaluate the effect of augmentations: The authors need to perform an ablation study to compare the different argumentation methods they proposed. 5. Analyze Training on Incorrect Answers: The authors should analyze whether wrong answers generated in data augmentation can harm the performance. 6. Evaluate other ways to increase the size of training data: The authors should evaluate other ways to increase the training data size and compare the performance with models trained on their proposed train data. 7. Error Analysis: The authors should analyze the performance of their models in different conditions (e.g., lengths of questions). Prompt Sensitivity Analysis: Evaluate the sensitivity of MetaMath to different prompt formats or phrasings of mathematical questions. B
Large Language Models Cannot Self-Correct Reasoning Yet 1. Self-Correction with Oracle Labels: The authors should evaluate self-correction performance with oracle labels. 2. Intrinsic Self-Correction: The authors should show performance without using the oracle labels. 3. Analysis of Mistakes in Self-Correction: The authors should analyze the properties of mistakes made in the self-correction framework. 4. Multi-Agent Debate: The authors should evaluate self-correction with multi-agent debate. 5. Prompt Design Analysis: The authors should analyze the influence of prompt design for the initial responses on self-correction performance. Visualization of learned representations or attention mechanisms to provide insights into the model’s inner workings. C
C.3 Human Evaluation on LLM-Generated Explanation
@@
We ask 5 annotators to evaluate the LLM-generated explanations. Specifically, each of them is assigned 4 or 5 papers, along with the corresponding experiment lists. For each paper, the annotator is given 5 different models’ outputs (model names are anonymized), and the annotator has to decide if each LLM-generated explanation is acceptable according to the experiment. We show the human evaluation results in Table 5.
C.4 Multi-Modal Input Ablation
We post the multi-modal ablation study of ExpDesign and Weakness in Table 14 and Table 15.
Experiment Design Experiment Explanation
Models En-F1 En-Precision En-Recall S-Match ROUGE-L ROUGE-1
GPT-4o 25.03 22.25 36.59 58.54 29.25 35.50
w/ figures 25.39 24.35 32.80 58.53 27.87 34.30
InternVL2-26B 24.26 39.50 14.91 50.03 29.13 34.26
w/ figures 15.04 38.50 8.64 50.29 29.29 34.06
Models S-F1 S-Precision S-Recall ITF-IDF
GPT-4o 47.73 42.09 55.48 5.95
w/ tables 46.76 41.32 54.17 5.53
w/ figures 46.62 41.20 54.04 5.48
w/ tables & figures 46.58 41.17 53.98 5.36
InternVL2-26B 41.91 41.02 43.28 1.48
w/ tables 40.55 40.37 42.91 1.46
w/ figures 42.88 42.10 43.76 1.46
w/ tables & figures 42.44 42.00 43.31 1.44